# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 4 Token-level Optimization
In the token level, optimization focuses exclusively on improving KV cache based on the characteristics and patterns of KV pairs of tokens, without considering enhancements from model architecture improvements or system parallelization techniques. In general, token-level optimization methods are primarily guided by observations from LLMs and sequential inputs. Existing approaches can be categorized into five main types: KV cache selection, KV cache budget allocation, KV cache merging, KV cache quantization, and KV cache low-rank decomposition. The taxonomy of the token-level optimization is shown in Fig. 3.
@@ 4.1 KV Cache Selection
KV cache selection mechanisms have emerged as a critical optimization strategy, aimed at reducing memory utilization of KV caches, minimizing inference latency, and enhancing overall throughput in large language models. These optimization objectives have driven the development of various selection methodologies, which can be classified into two distinct categories: (1) static KV cache selection, which performs token filtering exclusively during the prefilling phase, with selected tokens remaining fixed throughout subsequent decoding steps; and (2) dynamic KV cache selection, which continuously updates KV cache during the decoding phase, enabling adaptive cache management. In dynamic KV cache selection approaches, KV cache tokens that are not selected may be permanently evicted or offloaded to hierarchical caching devices such as CPU memory, implementing a multi-tier storage strategy. Given that real-time KV cache selection during decoding may incur substantial computational overhead, several studies have focused on developing optimized retrieval algorithms to enhance the efficiency of this process. These optimizations include block-level retrieval instead of token-level granularity to reduce search complexity, asynchronous query mechanisms to hide latency, and parallel retrieval pipelines to accelerate the selection process. These optimization efforts aim to mitigate the computational burden while maintaining the effectiveness of token selection. The summary of the KV cache selection is listed in Tab. II.
@@
Method Initial tokens Top-$k$ tokens Recent tokens Permanent eviction Dynamic selection Selection granularity Remark
@@ 4.1.1 Static KV Cache Selection
Static KV cache selection methods perform a one-time compression on the KV Cache immediately after the prefilling phase is completed. The model then uses this compressed KV cache for subsequent decoding inference. FastGen [133] introduces a pattern-aware approach by identifying five fundamental attention structures and implementing targeted selection strategies. These include proximity-based retention for local attention patterns, selective preservation of critical tokens for punctuation-focused attention, frequency-based filtering for sparse attention distributions, and complete token retention for broad attention patterns. SnapKV [134] simplifies FastGen’s approach by focusing solely on retrieving tokens based on their importance scores. It demonstrates that among all prompt tokens, only a portion carries crucial information for response generation, with these tokens maintaining their significance during the generation phase. The approach employs an end-positioned observation window to detect these important contextual tokens. Their corresponding key-value pairs are then concatenated with the tokens from the observation window. Attention-Gate [135] introduces a learnable KV-Cache eviction mechanism that processes the entire context sequence and generates token-wise eviction decisions through a parameterized policy network, enabling dynamic in-context memory management.
4.1.2 Dynamic Selection with Permanent Eviction
Recent works have explored leveraging attention scores as a criterion for selecting significant KV cache tokens. H2O [127] observes that attention computations are primarily driven by a select group of high-impact tokens, known as Heavy Hitters (H2). This method reformulates cache optimization as a dynamic submodular problem, utilizing cumulative attention scores to guide token retention decisions. Unlike H2O, BUZZ [128] employs a beehive-like structure that selects Heavy Hitters in local KV cache segments. NACL [129] identifies a fundamental limitation in H2O, namely their dependence on potentially biased local attention statistics. To overcome this issue, they develop an alternative approach implementing a diversified random eviction strategy for token selection. Scissorhands [130] builds upon the temporal significance principle, which suggests that tokens demonstrating historical importance maintain their influence in subsequent computational steps. This observation enables the preservation of repetitive attention patterns through selective token retention. Additionally, Keyformer [131] reveals that token removal distorts the underlying softmax probability distribution. Considering the pivotal role of softmax distributions in token significance evaluation, they incorporate regularization techniques to mitigate these distributional perturbations. SepLLM [132] observes that separator tokens (e.g., commas, periods, and line breaks) receive disproportionately high attention scores and naturally summarize text segments. Building on this, SepLLM retains separator tokens together with initial tokens, important tokens, and recent tokens in the cache.
4.1.3 Dynamic Selection without Permanent Eviction
To accelerate the retrieval of critical tokens, several research efforts have proposed index-based approaches that organize and access KV cache at block or cluster granularity, enabling efficient query and extraction operations. InfLLM [121] maintains full KV cache in blocks while facilitating long sequence processing through a hierarchical storage strategy. The framework employs CPU-GPU memory orchestration, preserving essential tokens and current computational units in GPU memory while offloading less frequently accessed units to CPU memory. To further enhance top-$k$ block retrieval precision, the Quest [122] framework presents a refined block representation approach based on minimal and maximal key values in KV cache blocks. PQCache [98] also implements block-based KV cache management and identifies salient tokens through Maximum Inner-Product Search (MIPS), leveraging Product Quantization (PQ) codes and centroids. SqueezedAttention [123] employs K-means clustering in an offline stage to group semantically similar keys, with each group represented by a centroid. During inference, it compares input queries against these centroids to identify and load only the semantically relevant keys from the context. Similarly, RetrievalAttention [124] index KV cache tokens using approximate nearest neighbor search (ANNS) techniques. Additionally, EM-LLM [125] dynamically segments incoming tokens into episodic events. Besides, it implements a hybrid retrieval mechanism that combines semantic similarity matching with temporal context to efficiently access relevant KV cache segments. Similarly, ClusterKV [126] groups tokens into semantic clusters and selectively recalls them during inference, achieving both high accuracy and efficiency for LLMs.
@@ 4.1.3 Dynamic Selection without Permanent Eviction
To accelerate top-$k$ token identification, SparQ [138] identifies the $r$ most significant elements in the incoming query vector and selectively retrieves the corresponding components along the hidden dimension of the cached key matrix $K$ for approximate attention computation. To overlap prefetching latency, InfiniGen [139] employs asynchronous prefetching, utilizing indices of salient KV entries selected by queries from the previous layer to retrieve KV cache entries in the current layer. To ensure maximum model performance, RecycledAttention [140] sustains the entire KV cache during inference computations, yielding no improvements in memory efficiency. The approach performs periodic top-$k$ token selection to identify salient tokens. Moreover, MagicPIG [141] shows that attention-based top-$k$ selection may incur performance degradation. To address this limitation, they introduce a novel heterogeneous computing framework leveraging Locality Sensitive Hashing (LSH) techniques. The system stores LSH hash tables and performs attention estimation on CPU.
@@ 4.1.4 Summary and Future Directions
Static KV cache selection algorithms demonstrate superior decoding efficiency overall; however, their efficacy remains to be thoroughly validated in multi-turn dialogues and extended decoding length scenarios. Dynamic KV cache selection algorithms, while adaptive, introduce additional computational overhead during the decoding phase due to frequent cache selection operations. Multi-tier cache architectures and prefetching schemes partially mitigate these challenges, yet their capability to achieve rapid and accurate retrieval within acceptable decoding latency constraints requires further empirical validation, particularly in real-world applications involving long sequences. Furthermore, existing selection methods predominantly rely on attention score-based top-$k$ selection mechanisms. However, based on existing positional encoding schemes, current top-$k$ approaches may not be able to effectively identify and extract relevant tokens in ultra-long sequence tasks.
@@ 4.2 KV Cache Budget Allocation
The hierarchical architecture of LLMs leads to diverse information extraction patterns across layers, with each layer’s KV-cache contributing differently to model performance. This inherent heterogeneity indicates that uniform KV-cache compression across layers may be suboptimal. KV cache budget allocation addresses this challenge by intelligently distributing memory resources based on each component’s importance to prediction accuracy, thereby optimizing memory utilization while minimizing accuracy degradation. Current budget allocation strategies can be categorized into two levels of granularity: layer-wise budget allocation, which assigns different compression ratios across model layers, and the more fine-grained head-wise budget allocation, which enables precise memory distribution across individual attention heads within each layer, offering more flexible and targeted optimization opportunities. The summary of KV budget allocation is listed in Tab. III.
@@
4.2.1 Layer-wise Budget Allocation
In contrast to conventional approaches with uniform KV cache sizes, PyramidKV [116] employs a pyramid-shaped memory allocation strategy, assigning larger cache capacities to lower layers that progressively decrease in upper layers. This design is supported by the observation that lower layers exhibit uniform attention distributions across input sequences, while upper layers show concentrated attention on specific tokens. PyramidInfer [117] also adopts a pyramid-shaped budget allocation strategy while selecting tokens with high attention values at each layer. Additionally, during the decoding phase, PyramidInfer dynamically maintains a set of significant tokens through frequent updates driven by attention values. Unlike previous methods, DynamicKV [118] implements an input-adaptive budget allocation strategy by analyzing attention patterns. Specifically, it computes the average attention scores between recent and historical tokens, identifies the top-$k$ tokens with highest attention values across layers, and proportionally distributes the budget based on the density of significant tokens in each layer. Similarly, PrefixKV [119] identifies the most important tokens for each layer by computing the average attention score of tokens within that layer. PrefixKV [119] then uses a unified threshold to determine the number of retained tokens, adaptively adjusting the retention for each layer based on its importance distribution. CAKE [142] examines attention scores through two lenses: the spatial distribution of inter-token attention and the temporal evolution of attention focus. These measurements are combined to compute layer-specific importance scores, which further guide the allocation of memory resources. Additionally, SimLayerKV [120] identifies lazy layers - those exhibiting limited effectiveness in capturing long-range dependencies. The framework then selectively preserves cache entries, maintaining initial and recent tokens for lazy layers while retaining complete KV cache for non-lazy layers.
4.2.2 Head-wise Budget Allocation
AdaKV [110] leverages the observation that attention patterns exhibit distinct concentrations across different heads. It implements head-specific memory allocation by optimizing an L1 loss bound between the original and pruned multi-head attention outputs. Within the constraints of a layer-wise budget, the method distributes cache capacity among heads to maximize the preserved attention information collectively. Building upon AdaKV, CriticalKV [111] introduces significant enhancements by recognizing that the importance of KV cache entries extends beyond attention weights to encompass value states and pretrained parameter matrices. Leveraging this insight, the framework implements a novel selection algorithm that identifies essential cache entries by minimizing the maximum potential output perturbation. LeanKV [112] implements a fine-grained memory optimization strategy that operates independently for each attention head and input request. The method identifies the smallest subset of tokens necessary to preserve the majority of information flow, allocating cache space based on a predefined attention score threshold - typically maintaining 95% of the total attention mass.
Retrieval head-based methods represent a specialized category of head-wise allocation strategies that focuses on identifying and prioritizing attention heads crucial for extracting key information from long sequences. This approach allocates larger cache budgets to these specialized heads, known as retrieval heads [143], due to their significant role in information extraction. RazorAttention [113] characterizes two distinct categories of retrieval heads: echo heads, which focus on previously occurring identical tokens, and induction heads, which attend to antecedent tokens that precede current token repetitions. This framework implements differential caching strategies, maintaining complete cache entries for retrieval heads while condensing remote tokens into consolidated compensation tokens for non-retrieval heads. HeadKV [114] further enhances RazorAttention by introducing a novel head assessment framework that simultaneously evaluates both retrieval and reasoning capabilities to optimize KV cache allocation strategies. DuoAttention [115] further introduces a parameterized approach to distinguish between two categories of attention mechanisms: retrieval heads, essential for comprehensive long-context processing, and Streaming heads, which primarily engage with recent tokens and attention sinks. This classification is achieved through learned parameters that automatically identify retrieval heads requiring full attention spans.
4.2.3 Summary and Future Directions
Despite recent advances and growing attention in KV cache budget allocation research, several critical challenges remain unaddressed. First, the relationship between allocation strategies and model performance requires further investigation. For instance, a notable discrepancy exists between pyramid-shaped allocation strategies [116, 117] advocating larger budgets for lower layers, and retrieval head-based studies [113, 114] which demonstrate that lower layers rarely exhibit retrieval head characteristics and thus require minimal cache resources. Additionally, the field lacks comprehensive experimental comparisons, particularly regarding the compatibility and performance benefits of head-wise budget allocation strategies with state-of-the-art frameworks like vLLM [144] and FlashAttention [145]. Also, existing methods, such as PyramidInfer [117], demonstrate some adaptability to input attention patterns. However, future research could target real-time, task-specific allocation strategies that dynamically adjust memory budgets during inference based on input characteristics, task complexity, or downstream requirements.
@@ 4.3.1 Intra-layer Merging
Instead, CaM [104], KVMerger [108], and D2O [105] are training-free, which rely on observations and directly propose rule-based or heuristic-based merging strategies. Specifically, they separate the Keys and Values of tokens in each layer into important (retrained) and unimportant (evicted) tokens. They then keep potentially useful unimportant tokens by merging their Keys and Values with retained important tokens, ensuring that no valuable information is lost. Particularly, D2O [105] merges merges the Key (or Value) of a evicted token with one retained token based on cosine similarity. Similar to D2O based on cosine similarity, AIM [106] and Look-M [107] merges Keys (resp. Values) of multiple tokens into one. CaM [104] merges the Keys (or Values) of multiple evicted tokens with retained tokens based on attention scores to get the final merged results. Also, KVMerger [108] first identifies the merge token sets by clustering consecutive tokens with high cosine similarity, ensuring that only adjacent tokens with strong contextual relevance are grouped together. Then, KVMerger merges the tokens in each merge set into the pivotal token (chosen based on the highest attention score) using Gaussian kernel weights, where closer tokens contribute more to the merged state.
@@ 4.4 KV Cache Quantization
Quantization technique [151, 152, 153, 154, 155] has been widely used to accelerate machine learning models from different aspects, such model parameter quantization [156, 157, 158, 159] and data feature quantization [160, 161]. Similarly, Key-Value (KV) cache quantization is emerging as a highly promising solution to address the memory and computational bottlenecks in LLMs. During autoregressive decoding, LLMs generate key-value pairs for every attention layer across all tokens in the sequence. If we store all KV pairs in the memory with full precision, this cache grows exponentially with longer sequences, increasing the memory and bandwidth requirements significantly. Quantization reduces the precision of numerical representations (e.g., from FP32 to INT8 or INT4), drastically compressing the size of the KV cache. This compression can achieve up to 4x or more memory savings, making it feasible for LLMs to operate on resource-constrained devices like GPUs with limited memory or edge devices.
@@ 4.4.1 Fixed-precision Quantization
Fixed-precision quantization proposes quantizing different Keys (different Values) of tokens to the same bit-width. ZeroQuant [95] propose per-token quantization for Keys and Values. As shown in Fig. 4, the per-token quantization approach quantizes tokens individually. Particularly, ZeroQuant [95] dynamically computes the min-max range for each token during inference. This ensures that each token is quantized based on its unique range, significantly reducing quantization error. Also FlexGen [96] and QJL [97] directly perform per-token quantization for Keys and Values, where the scaling factor and zero-point are shared among all elements within the same token. PQCache [98] uses product quantization approaches [161, 165] to compress KV pairs. However, uniform quantization approaches, which use a fixed bit-width for keys and values across all tokens, can often be suboptimal. It is because they ignore the varying importance of tokens [112] and account for the outlier patterns in Keys and Values [148, 84].

### v3

@@ 4 Token-level Optimization
At the token level, optimization focuses exclusively on improving the KV cache management based on the characteristics and patterns of the KV pairs of tokens, without considering enhancements from model architecture improvements or system parallelization techniques. Generally, token-level optimization methods are guided by observations from LLMs and sequential inputs. Existing approaches can be categorized into five main types: KV cache selection, KV cache budget allocation, KV cache merging, KV cache quantization, and KV cache low-rank decomposition. The taxonomy of token-level optimization is shown in Fig. 3.
@@ 4.1 KV Cache Selection
Figure 4: The sparsity of attention matrix.
As shown in Fig. 4, the attention matrix is sparse. KV cache selection mechanisms have emerged as a critical optimization strategy, aimed at reducing memory utilization of KV caches, minimizing inference latency, and enhancing overall throughput in large language models. These optimization objectives have driven the development of various selection methodologies, which can be classified into two distinct categories: (1) static KV cache selection, which performs token filtering exclusively during the prefilling phase, with selected tokens remaining fixed throughout subsequent decoding steps; and (2) dynamic KV cache selection, which continuously updates KV cache during the decoding phase, enabling adaptive cache management. In dynamic KV cache selection approaches, KV cache tokens that are not selected may be permanently evicted or offloaded to hierarchical caching devices such as CPU memory, implementing a multi-tier storage strategy. Given that real-time KV cache selection during decoding may incur substantial computational overhead, several studies have focused on developing optimized retrieval algorithms to enhance the efficiency of this process. These optimizations include block-level retrieval instead of token-level granularity to reduce search complexity, asynchronous query mechanisms to hide latency, and parallel retrieval pipelines to accelerate the selection process. These optimization efforts aim to mitigate the computational burden while maintaining the effectiveness of token selection. The summary of the KV cache selection is listed in Tab. II.
@@
[…]
@@ 4.1.1 Static KV Cache Selection
Static KV cache selection methods perform a one-time compression on the KV Cache immediately after the prefilling phase is completed. The model then uses this compressed KV cache for subsequent decoding inference. FastGen [137] introduces a pattern-aware approach by identifying five fundamental attention structures and implementing targeted selection strategies. These include proximity-based retention for local attention patterns, selective preservation of critical tokens for punctuation-focused attention, frequency-based filtering for sparse attention distributions, and complete token retention for broad attention patterns. SnapKV [138] simplifies FastGen’s approach by focusing solely on retrieving tokens based on their importance scores. It demonstrates that among all prompt tokens, only a portion carries crucial information for response generation, with these tokens maintaining their significance during the generation phase. The approach employs an end-positioned observation window to detect these important contextual tokens. Their corresponding key-value pairs are then concatenated with the tokens from the observation window. L2Compress [139] proposes another simple yet effective approach to identify important tokens during the prefilling phase using the $L_{2}$ norm for key embeddings. By analyzing attention distributions in decoder-only Transformers, the authors find a clear correlation between the $L_{2}$ norm of a key embedding and its attention score during decoding, where a low $L_{2}$ norm usually leads to high attention scores. Based on this observation, they propose a strategy that compresses the KV Cache by retaining tokens with lower $L_{2}$ norms. Attention-Gate [140] introduces a learnable KV-Cache eviction mechanism that processes the entire context sequence and generates token-wise eviction decisions through a parameterized policy network, enabling dynamic in-context memory management.
@@
4.1.2 Dynamic Selection with Permanent Eviction
Recent works have explored leveraging attention scores as a criterion for selecting significant KV cache tokens. H2O [131] observes that attention computations are primarily driven by a select group of high-impact tokens, known as Heavy Hitters. This method reformulates cache optimization as a dynamic submodular problem, utilizing cumulative attention scores to guide token retention decisions. Unlike H2O, BUZZ [132] employs a beehive-like structure that selects Heavy Hitters in local KV cache segments. NACL [133] identifies a fundamental limitation in H2O, namely their dependence on potentially biased local attention statistics. To overcome this issue, they develop an alternative approach implementing a diversified random eviction strategy for token selection. Scissorhands [134] builds upon the temporal significance principle, which suggests that tokens demonstrating historical importance maintain their influence in subsequent computational steps. This observation enables the preservation of repetitive attention patterns through selective token retention. Additionally, Keyformer [135] reveals that token removal distorts the underlying softmax probability distribution. Considering the pivotal role of softmax distributions in token significance evaluation, they incorporate regularization techniques to mitigate these distributional perturbations. SepLLM [136] observes that separator tokens (e.g., commas, periods, and line breaks) receive disproportionately high attention scores and naturally summarize text segments. Building on this, SepLLM retains separator tokens together with initial tokens, important tokens, and recent tokens in the cache.
4.1.3 Dynamic Selection without Permanent Eviction
[…]
@@ 4.1.4 Summary and Future Directions
[…]
@@ 4.2 KV Cache Budget Allocation
The hierarchical architecture of LLMs leads to diverse information extraction patterns across layers, with each layer’s KV-cache contributing differently to model performance. This inherent heterogeneity indicates that uniform KV-cache compression across layers may be suboptimal. KV cache budget allocation addresses this challenge by intelligently distributing memory resources based on each component’s importance to prediction accuracy, thereby optimizing memory utilization while minimizing accuracy degradation. It is worth noting that budget allocation approaches prioritize the effective allocation of computational budget rather than token selection strategies. Current budget allocation strategies can be categorized into two levels of granularity: layer-wise budget allocation, which assigns different compression ratios across model layers, and the more fine-grained head-wise budget allocation, which enables precise memory distribution across individual attention heads within each layer, offering more flexible and targeted optimization opportunities. The summary of KV budget allocation is listed in Tab. III.
4.2.1 Layer-wise Budget Allocation
[…]
4.2.2 Head-wise Budget Allocation
AdaKV [114] leverages the observation that attention patterns exhibit distinct concentrations across different heads. It implements head-specific memory allocation by optimizing an L1 loss bound between the original and pruned multi-head attention outputs. Within the constraints of a layer-wise budget, the method distributes cache capacity among heads to maximize the preserved attention information collectively. Building upon AdaKV, CriticalKV [115] introduces significant enhancements by recognizing that the importance of KV cache entries extends beyond attention weights to encompass value states and pretrained parameter matrices. Leveraging this insight, the framework implements a novel selection algorithm that identifies essential cache entries by minimizing the maximum potential output perturbation. LeanKV [116] implements a fine-grained memory optimization strategy that operates independently for each attention head and input request. The method identifies the smallest subset of tokens necessary to preserve the majority of information flow, allocating cache space based on a predefined attention score threshold and typically maintaining 95% of the total attention mass.
Retrieval head-based methods represent a specialized category of head-wise allocation strategies that focus on identifying and prioritizing attention heads crucial for extracting key information from long sequences. This approach allocates larger cache budgets to these specialized heads, known as retrieval heads [148], due to their significant role in information extraction. RazorAttention [117] characterizes two distinct categories of retrieval heads: echo heads, which focus on previously occurring identical tokens, and induction heads, which attend to antecedent tokens that precede current token repetitions. This framework implements differential caching strategies, maintaining complete cache entries for retrieval heads while condensing remote tokens into consolidated compensation tokens for non-retrieval heads. HeadKV [118] further enhances RazorAttention by introducing a novel head assessment framework that simultaneously evaluates both retrieval and reasoning capabilities to optimize KV cache allocation strategies. DuoAttention [119] further introduces a parameterized approach to distinguish between two categories of attention mechanisms: retrieval heads, essential for comprehensive long-context processing, and Streaming heads, which primarily engage with recent tokens and attention sinks. This classification is achieved through learned parameters that automatically identify retrieval heads requiring full attention spans.
4.2.3 Summary and Future Directions
Despite recent advances and growing attention in KV cache budget allocation research, several critical challenges remain unaddressed. First, the relationship between allocation strategies and model performance requires further investigation. For instance, a notable discrepancy exists between pyramid-shaped allocation strategies [120, 121] advocating larger budgets for lower layers, and retrieval head-based studies [117, 118] which demonstrate that lower layers rarely exhibit retrieval head characteristics and thus require minimal cache resources. Additionally, the field lacks comprehensive experimental comparisons, particularly regarding the compatibility and performance benefits of head-wise budget allocation strategies with state-of-the-art frameworks like vLLM [149] and FlashAttention [150] [151] [152]. Also, existing methods, such as PyramidInfer [121], demonstrate some adaptability to input attention patterns. However, future research could target real-time, task-specific allocation strategies that dynamically adjust memory budgets during inference based on input characteristics, task complexity, or downstream requirements.
@@ 4.3.1 Intra-layer Merging
Instead, CaM [107], KVMerger [112], ZeroMerge [155], and D2O [108] are training-free, which rely on observations and directly propose rule-based or heuristic-based merging strategies. Specifically, they separate the Keys and Values of tokens in each layer into important (retained) and unimportant (evicted) tokens. They then keep potentially useful unimportant tokens by merging their Keys and Values with retained important tokens, ensuring that no valuable information is lost. Particularly, D2O [108] merges the Key (or Value) of a evicted token with one retained token based on cosine similarity. Similar to D2O, AIM [109], Look-M [110], and ZeroMerge [111] merge Keys (resp. Values) of multiple tokens into one. CaM [107] merges the Keys (or Values) of multiple evicted tokens with retained tokens based on attention scores to get the final merged results. Also, KVMerger [112] first identifies the merge token sets by clustering consecutive tokens with high cosine similarity, ensuring that only adjacent tokens with strong contextual relevance are grouped together. Then, KVMerger merges the tokens in each merge set into the pivotal token (chosen based on the highest attention score) using Gaussian kernel weights, where closer tokens contribute more to the merged state.
@@ 4.4 KV Cache Quantization
Figure 5: The quantization of attention matrix.
As shown in Fig.5, quantization techniques [159, 160, 161, 162, 163] aim to convert full-precision values into lower-bit integers, reducing computational and storage requirements. Quantization techniques have been widely used to accelerate machine learning models from different aspects, such model parameter quantization [164, 165, 166, 167] and data feature quantization [168, 169]. Similarly, Key-Value (KV) cache quantization is emerging as a highly promising solution to address the memory and computational bottlenecks in LLMs. During auto-regressive decoding, LLMs generate key-value pairs for every attention layer across all tokens in the sequence. If we store all KV pairs in the memory with full precision, this cache grows linearly with longer sequences, increasing the memory and bandwidth requirements significantly. Quantization reduces the precision of numerical representations (e.g., from FP32 to INT8 or INT4), drastically compressing the size of the KV cache. This compression can achieve up to 4x or more memory savings, making it feasible for LLMs to operate on resource-constrained devices like GPUs with limited memory or edge devices.
@@ 4.4.1 Fixed-precision Quantization
Fixed-precision quantization proposes quantizing different Keys (different Values) of tokens to the same bit-width. ZeroQuant [98] proposes per-token quantization for Keys and Values. As shown in Fig. 6, the per-token quantization approach quantizes tokens individually. Particularly, ZeroQuant [98] dynamically computes the min-max range for each token during inference. This ensures that each token is quantized based on its unique range, significantly reducing quantization error. Also FlexGen [99] and QJL [100] directly perform per-token quantization for Keys and Values, where the scaling factor and zero-point are shared among all elements within the same token. PQCache [101] uses product quantization approaches [169, 173] to compress KV pairs. However, uniform quantization approaches, which use a fixed bit-width for keys and values across all tokens, can often be suboptimal. It is because they ignore the varying importance of tokens [116] and account for the outlier patterns in Keys and Values [156, 87].